Please help me solve this problem.
A total of 508 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was three times the number of adult tickets sold. How many adult tickets were sold?

Answer by stanbon(75887)

Please help me solve this problem.
A total of 508 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was three times the number of adult tickets sold. How many adult tickets were sold?
---
Quantity Eq: a + s = 508
Quantity Eq: s = 3a
-----
Substitute for "s" and solve for "a":
a + 3a = 508
4a = 508
a = 127 (# of adult tickets sold)
------
Solve for "s":
s = 3a
s = 3*127
s = 381 (# of student tickets sold)
=====================================
Cheers,
Stan H.
===============